# The Complexity of One or Many Faces in the Overlay of Many Arrangements

### v1

@@ 1 Introduction
Let $\Gamma$ be a given set of $n$ Jordan arcsfootnoteA Jordan arc is the image of a continuous 111\!\!-\!\!1 mapping of the unit interval $[0,1]$ into the plane.in the plane, so that any pair of arcs from $\Gamma$ intersect at most $s$ times (for some fixed constant $s$). Let $A(\Gamma)$ denote the arrangement of $\Gamma$, namely the partition of the plane induced by the arcs of $\Gamma$ into $O(n^{2})$ faces, edges and vertices. The vertices are the endpoints and the points of intersection of the arcs in $\Gamma$, the edges are the maximal connected portions of the arcs not containing any vertex, and the faces are the connected components of the complement of the union of the arcs in $\Gamma$. Such an arrangement is depicted in Figure 1.1. See [SA95] for more details concerning planar arrangements of arcs.
@@ 2 The Complexity of a Single Face in an Overlay of Arrangements
In this section we derive a tight bound on the maximum complexity of a single face in an overlay of many arrangements of arcs. This bound is used in Section 3.1 to prove our main theorem. We also derive several related results, which we believe to be interesting in their own right.
@@ 2.3 Restricted Davenport-Schinzel Sequences
In this section we prove a bound on the length of a Davenport-Schinzel sequence, expressed in terms of the structure of the sequence restricted to subsets of its symbols. If $S$ is a sequence and $A$ is a subset of its symbols, we denote by $S_{A}$ the subsequence of $S$ obtained by removing all elements not in $A$, and by replacing maximal runs of adjacent equal elements by a single element.
@@
Proof of Lemma 2.2 Let $S$ be the lower envelope sequence of $F$, and let $S_{1},\ldots,S_{k}$ denote the lower envelope sequences of $F_{1},\ldots,F_{k}$, respectively. It can be easily checked that $S_{F_{i}}$ is a subsequence of $S_{i}$, for $i=1,\ldots,k$. Thus the total length of the restricted sequences of $S$ over $F_{1},\ldots,F_{k}$ is bounded by $C$.
@@
$|S|=O\left({C\frac{\lambda_{s}(k)}{k}}\right)$ (or, for partially-defined functions, $|S|=O\left({C\frac{\lambda_{s+2}(k)}{k}}\right))$, as claimed.
Definition Let $S=(s_{1},\ldots,s_{r})$ be a $DS(n,s)$ sequence. A symbol $a$ is active in position $j$ in $S$ if there are two indices $1\leq j^{\prime}\leq j<j^{{\prime\prime}}\leq r$ such that $s_{j^{\prime}}=s_{j^{{\prime\prime}}}=a$. Let $v(j)$ denote the number of active symbols in position $j$ of $S$. We call a sequence $S$ a $DSA(n,s,k)$ sequence if $S$ is a $DS(n,s)$ sequence and, for each $1\leq j\leq|S|$, we have $v(j)\leq k$. Let $\lambda_{s,k}(n)$ denote the maximum length of a $DSA(n,s,k)$ sequence.
@@ 2.4 The Complexity of a Single Face in an Overlay of Arrangements
In this section we finally prove the upper bound of Theorem 2.1.
@@
Figure 2.2: Traversing a face boundary and the resulting sequence $S=\left\{\gamma_{1}^{+}\right$. $\gamma_{2}^{-}$ $\gamma_{2}^{+}$ $\gamma_{1}^{+}$ $\gamma_{7}^{-}\gamma_{3}^{-}\gamma_{6}^{+}\gamma_{6}^{-}\gamma_{3}^{-}\gamma_{5}^{+}\gamma_{5}^{-}\gamma_{3}^{-}\gamma_{4}^{+}\gamma_{2}^{-}\left.\gamma_{1}^{-}\right\}$.
@@
because no arc of any sub-arrangement can appear along two distinct components of $\partial{f}$. This completes the proof of Theorem 2.1.
@@ 2.5 The Complexity of a Face in Certain Arrangements of Line Segments
In this section we present new bounds, for some special cases, on the complexity of a single face in an arrangement of line segments. The following lemma is the main tool in proving these new results.
@@
Definition Given a set $S$ of $n$ Jordan arcs, we define the covering number of $S$ to be the minimal size of a subset $T\subseteq S$, such that each arc of $S$ intersects at least one arc of $T$.
@@ 3.1 Many Faces in the Overlay of Many Arrangements
Substituting the bound of Lemma 3.3 into Lemma 3.2, we obtain a more explicit version of Lemma 3.2:
@@ 3.2 The Complexity of Many Faces in Sparse Arrangements
In this section we derive some bounds on the complexity of many faces in sparse arrangements, namely arrangements whose complexity is much smaller than quadratic (in the number of arcs).
Definition Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{n}\}$ be a set of $n$ Jordan arcs in the plane, so that each pair of arcs from $\Gamma$ have at most $s$ intersection points. We call a function $f:\Gamma\rightarrow\{1,\ldots,t\}$ a coloring of $\Gamma$ by $t$ colors, if $f(\gamma_{i})\neq f(\gamma_{j})$ for each intersecting pair of arcs $\gamma_{i},\gamma_{j}\in\Gamma$. The smallest number $t$, for which there exists a coloring of $\Gamma$ with $t$ colors, is called the chromatic number of $\Gamma$, and is denoted by $\chi(\Gamma)$.
@@
In the following analysis, we will use Theorem 3.6 of Clarkson and Shor [CS89], which shows that the expected value of expressions of the form $\sum_{i=1}^{\tau}W\left({n_{i}\choose d}\right)$, where $W$ is an arbitrary concave non-negative function and $d$ is a positive integer, satisfies
$E\left({\sum_{i=1}^{\tau}W\left({{{n_{i}}\choose{d}}}\right)}\right)\leq E(\tau)\cdot W\left({D\left({\frac{n}{r}}\right)^{d}}\right),$
@@
$E\left({\sum_{i=1}^{\tau}\sqrt{m_{i}}n_{i}}\right)$ $\leq$ $E\left({\sqrt{\sum_{i=1}^{\tau}m_{i}}\sqrt{\sum_{i=1}^{\tau}n_{i}^{2}}}\right)=\sqrt{m}\cdot E\left({\sqrt{\sum_{i=1}^{\tau}n_{i}^{2}}}\right)$
$\leq$ $\sqrt{m}\cdot\sqrt{E\left({\sum_{i=1}^{\tau}n_{i}^{2}}\right)},$
where the last inequality follows from the inequality $E\left({\sqrt{X}}\right)\leq\sqrt{E\left({X}\right)}$. Since $n_{i}^{2}=2{{n_{i}}\choose{2}}+n_{i}$, we obtain, using Theorem 3.6 of [CS89] with $W(x)\equiv x^{2}$, that
$E\left({\sum_{i=1}^{\tau}\sqrt{m_{i}}n_{i}}\right)$ $\leq$ $\sqrt{m}\sqrt{2E\left({\sum_{i=1}^{\tau}{{n_{i}}\choose{2}}}\right)+E\left({\sum_{i=1}^{\tau}n_{i}}\right)}$
@@
Acknowledgment
I wish to thank Micha Sharir, my thesis advisor, for his help in preparing the paper. He has also contributed to the proof of Theorem 3.12.

### v2

@@ 1 Introduction
Let $\Gamma$ be a given set of $n$ Jordan arcs [^2] A Jordan arc is the image of a continuous 111\!\!-\!\!1 mapping of the unit interval $[0,1]$ into the plane.in the plane, so that any pair of arcs from $\Gamma$ intersect at most $s$ times (for some fixed constant $s$). Let $A(\Gamma)$ denote the arrangement of $\Gamma$, namely the partition of the plane induced by the arcs of $\Gamma$ into $O(n^{2})$ faces, edges and vertices. The vertices are the endpoints and the points of intersection of the arcs in $\Gamma$, the edges are the maximal connected portions of the arcs not containing any vertex, and the faces are the connected components of the complement of the union of the arcs in $\Gamma$. Such an arrangement is depicted in Figure 1.1. See [SA95] for more details concerning planar arrangements of arcs.
@@ 2 The Complexity of a Single Face in an Overlay of Arrangements
In this section, we derive a tight bound on the maximum complexity of a single face in an overlay of many arrangements of arcs. This bound is used in Section 3.1 to prove our main theorem. We also derive several related results, which we believe to be interesting in their own right.
@@ 2.3 Restricted Davenport-Schinzel Sequences
In this section, we prove a bound on the length of a Davenport-Schinzel sequence, expressed in terms of the structure of the sequence restricted to subsets of its symbols. If $S$ is a sequence and $A$ is a subset of its symbols, we denote by $S_{A}$ the subsequence of $S$ obtained by removing all elements not in $A$, and by replacing maximal runs of adjacent equal elements by a single element.
@@
Proof of
Lemma 2.2: Let $S$ be the lower envelope sequence of $F$, and let $S_{1},\ldots,S_{k}$ denote the lower envelope sequences of $F_{1},\ldots,F_{k}$, respectively. It can be easily checked that $S_{F_{i}}$ is a subsequence of $S_{i}$, for $i=1,\ldots,k$. Thus the total length of the restricted sequences of $S$ over $F_{1},\ldots,F_{k}$ is bounded by $C$.
@@
$|S|=O\left({C\frac{\lambda_{s}(k)}{k}}\right)$ (or, for partially-defined functions, $|S|=O\left({C\frac{\lambda_{s+2}(k)}{k}}\right))$, as claimed. [Uncaptioned image]
Definition 2.5.
Let $S=(s_{1},\ldots,s_{r})$ be a $DS(n,s)$ sequence. A symbol $a$ is active in position $j$ in $S$ if there are two indices $1\leq j^{\prime}\leq j<j^{{\prime\prime}}\leq r$ such that $s_{j^{\prime}}=s_{j^{{\prime\prime}}}=a$. Let $v(j)$ denote the number of active symbols in position $j$ of $S$. We call a sequence $S$ a $DSA(n,s,k)$ sequence if $S$ is a $DS(n,s)$ sequence and, for each $1\leq j\leq|S|$, we have $v(j)\leq k$. Let $\lambda_{s,k}(n)$ denote the maximum length of a $DSA(n,s,k)$ sequence.
@@ 2.4 The Complexity of a Single Face in an Overlay of Arrangements
In this section, we finally prove the upper bound of Theorem 2.1.
@@
Figure 2.2: Traversing a face boundary and the resulting sequence $S=\{\gamma_{1}^{+}$$\gamma_{2}^{-}$$\gamma_{2}^{+}$$\gamma_{1}^{+}$$\gamma_{7}^{-}$$\gamma_{3}^{-}$$\gamma_{6}^{+}$$\gamma_{6}^{-}$$\gamma_{3}^{-}$$\gamma_{5}^{+}$$\gamma_{5}^{-}$$\gamma_{3}^{-}$$\gamma_{4}^{+}\gamma_{2}^{-}\gamma_{1}^{-}\}$.
@@
because no arc of any sub-arrangement can appear along two distinct components of $\partial{f}$. This completes the proof of Theorem 2.1. [Uncaptioned image]
@@ 2.5 The Complexity of a Face in Certain Arrangements of Line Segments
In this section, we present new bounds, for some special cases, on the complexity of a single face in an arrangement of line segments. The following lemma is the main tool in proving these new results.
@@
Definition 2.19.
Given a set $S$ of $n$ Jordan arcs, we define the covering number of $S$ to be the minimal size of a subset $T\subseteq S$, such that each arc of $S$ intersects at least one arc of $T$.
@@ 3.1 Many Faces in the Overlay of Many Arrangements
Substituting the bound of Lemma 3.3 into Lemma 3.2, we obtain the following more explicit version of Lemma 3.2:
@@ 3.2 The Complexity of Many Faces in Sparse Arrangements
In this section, we derive some bounds on the complexity of many faces in sparse arrangements, namely arrangements whose complexity is much smaller than quadratic (in the number of arcs).
Definition 3.7.
Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{n}\}$ be a set of $n$ Jordan arcs in the plane, so that each pair of arcs from $\Gamma$ have at most $s$ intersection points. We call a function $f:\Gamma\rightarrow\{1,\ldots,t\}$ a coloring of $\Gamma$ by $t$ colors, if $f(\gamma_{i})\neq f(\gamma_{j})$ for each intersecting pair of arcs $\gamma_{i},\gamma_{j}\in\Gamma$. The smallest number $t$, for which there exists a coloring of $\Gamma$ with $t$ colors, is called the chromatic number of $\Gamma$, and is denoted by $\chi(\Gamma)$.
@@
In the following analysis, we will use Theorem 3.6 of Clarkson and Shor [CS89], which shows that the expected value of expressions of the form $\sum_{i=1}^{\tau}W\left(\binom{n_{i}}{d}\right)$, where $W$ is an arbitrary concave non-negative function and $d$ is a positive integer, satisfies
$E\left({\sum_{i=1}^{\tau}W\left({\binom{n_{i}}{d}}\right)}\right)\leq E(\tau)\cdot W\left({D\left({\frac{n}{r}}\right)^{d}}\right),$
@@
$E\left({\sum_{i=1}^{\tau}\sqrt{m_{i}}n_{i}}\right)$ $\leq$ $E\left({\sqrt{\sum_{i=1}^{\tau}m_{i}}\sqrt{\sum_{i=1}^{\tau}n_{i}^{2}}}\right)=\sqrt{m}\cdot E\left({\sqrt{\sum_{i=1}^{\tau}n_{i}^{2}}}\right)\leq\sqrt{m}\cdot\sqrt{E\left({\sum_{i=1}^{\tau}n_{i}^{2}}\right)},$
where the last inequality follows from the inequality $E\left({\sqrt{X}}\right)\leq\sqrt{E\left({X}\right)}$. Since $n_{i}^{2}=2{\binom{n_{i}}{2}}+n_{i}$, we obtain, using Theorem 3.6 of [CS89] with $W(x)\equiv x^{2}$, that
$E\left({\sum_{i=1}^{\tau}\sqrt{m_{i}}n_{i}}\right)$ $\leq$ $\sqrt{m}\sqrt{2E\left({\sum_{i=1}^{\tau}{\binom{n_{i}}{2}}}\right)+E\left({\sum_{i=1}^{\tau}n_{i}}\right)}$
@@
Acknowledgments.
I wish to thank Micha Sharir, my thesis advisor, for his help in preparing the paper. He has also contributed to the proof of Theorem 3.13. My thanks also go to David Zheng for spotting typos in the arXiv version—and for proving me wrong in my assumption that no one would ever look at it.